OpenStudy (hanner_b_nanner):

Fan&&Medal I have no idea how to do this Simplify the quantity x plus 6 over 12, all over the quantity x minus 8 over 10 A 5 x over quantity x minus 8 B 5 x over 2 open parentheses x minus 4 close parentheses C 5 open parentheses x plus 6 close parentheses over 6 open parentheses x minus 8 close parentheses D open parentheses x plus 6 close parentheses open parentheses x minus 8 close parentheses over 30

OpenStudy (campbell_st):

if you are dividing by a fraction, you need to flip the denominator and multiply so it becomes \[\frac{x + 6}{12} \times \frac{10}{x + 8}\]

OpenStudy (campbell_st):

this can be simplified by identifying and deviding by the common factor

OpenStudy (hanner_b_nanner):

ok, everything can be divided by 2.. so would it change to x+3 over 6 times 5 over x+4?

OpenStudy (campbell_st):

no, just divide the 10 and 12 by 2. Then (x + 6) and (x + 8) are unchanged so 5(x + 6)/6(x + 4)

OpenStudy (campbell_st):

oops 5(x + 6)/6(x + 8)

OpenStudy (hanner_b_nanner):

ohhhhh

OpenStudy (hanner_b_nanner):

Thank you so much!